OpenStudy (anonymous):

Find the equation of the tangent line of the given function at the given point. g(x) =1/square root x , at (1,1)

OpenStudy (freckles):

I seen earlier you haven't gotten to short cuts But you should know the definition of derivative and that is what I bet you are to use: \[g'(x)=\lim_{h \rightarrow 0}\frac{g(x+h)-g(x)}{h}\]

OpenStudy (freckles):

once we find g'(x) we can find g'(1) to find the slope of our tangent line

OpenStudy (freckles):

the equation of the tangent line at (1,g(1)) will be in the following form: \[y-g(1)=g'(1)(x-1)\]

OpenStudy (freckles):

g(1) is actually already given as 1 you real task is to really just find g'(1) and replace it in that equation I gave

OpenStudy (anonymous):

\[\lim_{x \rightarrow 1}\frac{ 1 }{ \sqrt{x} }-1 /(x-1)\] ????

OpenStudy (freckles):

hmm.. did you try to use the definition of derivative I gave you?

OpenStudy (freckles):

\[g'(x)=\lim_{h \rightarrow 0}\frac{g(x+h)-g(x)}{h} \\ =\lim_{h \rightarrow 0}\frac{\frac{1}{\sqrt{x+h}}-\frac{1}{\sqrt{x}}}{h}=\lim_{h \rightarrow 0}\frac{1}{h}(\frac{1}{\sqrt{x+h}}-\frac{1}{\sqrt{x}})\] combine the fractions in the ( ) and then you will need to rationalize the numerator after that

OpenStudy (anonymous):

okay, let me try that..

OpenStudy (anonymous):

how do i get rid of the square roots?

OpenStudy (freckles):

rationalizing (sqrt(a)-sqrt(b))(sqrt(a)+sqrt(b)) =a-b

OpenStudy (anonymous):

can't do it :( i got 1/x^2...

OpenStudy (freckles):

so did you combine the fractions inside the ( )

OpenStudy (anonymous):

i think so...

OpenStudy (freckles):

\[g'(x)=\lim_{h \rightarrow 0}\frac{g(x+h)-g(x)}{h} \\ =\lim_{h \rightarrow 0}\frac{\frac{1}{\sqrt{x+h}}-\frac{1}{\sqrt{x}}}{h}=\lim_{h \rightarrow 0}\frac{1}{h}(\frac{1}{\sqrt{x+h}}-\frac{1}{\sqrt{x}}) \\ =\lim_{h \rightarrow 0}\frac{1}{h}\frac{\sqrt{x}-\sqrt{x+h}}{\sqrt{x+h} \sqrt{x}}\]

OpenStudy (freckles):

now rationalize the numerator

OpenStudy (freckles):

keep in mind whatever you multiply on top you multiply on bottom

OpenStudy (anonymous):

do i need to change the conjugate when rationalizing?

OpenStudy (freckles):

multiply top and bottom by the conjugate of the top

OpenStudy (anonymous):

 am i on the right track?